A bomb calorimeter, or a constant volume calorimeter, is a device often used to determine the heat of combustion of fuels and the energy content of foods. In an experiment, a 0.4395 g sample of biphenyl (C12H10) is burned completely in a bomb calorimeter. The calorimeter is surrounded by 1.064×103 g of water. During the combustion the temperature increases from 26.33 to 29.52 °C. The heat capacity of water is 4.184 J g-1°C-1. The heat capacity of the calorimeter was determined in a previous experiment to be 946.6 J/°C. Assuming that no energy is lost to the surroundings, calculate the molar heat of combustion of biphenyl based on these data. C12H10(s) + (29/2) O2(g) 5 H2O(l) + 12 CO2(g) + Energy Molar Heat of Combustion = kJ/mol
Answer #1

heat of combustion = heat capacity of calorimter x ∆T
Q = 946.6 J/ºC x 3.19ºC = 3019.654 J

moles of biphenyl = (0.4395g / 154g/mole) = 0.0028538 moles

molar heat capacity of biphenyl = (3019.654 J / 0.0028538 moles )= 1058081 J/mole = 1058.081 KJ/mol